### v1

# Enhancing Clinical Trial Patient Matching through Knowledge Augmentation with Multi-Agents

Abstract
Matching patients effectively and efficiently for clinical trials is a significant challenge due to the complexity and variability of patient profiles and trial criteria. This paper presents a novel framework, Multi-Agents for Knowledge Augmentation (MAKA), designed to enhance patient-trial matching by dynamically supplementing matching prompts with external, domain-specific knowledge. The MAKA architecture consists of five key components: a knowledge probing agent that detects gaps in domain knowledge, a navigation agent that manages interactions among multiple specialized knowledge augmentation agents, a knowledge augmentation agent that incorporates relevant information into patient-trial matching prompts, a supervision agent aligning the outputs from other agents with the instructions and a matching agent making the final selection decision. This approach enhances the accuracy and contextual richness of patient matching, addresses inherent knowledge gaps in both trail criteria and large language models (LLMs), and improves the alignment between patient characteristics and the criteria.
1 Introduction
Patient matching for clinical trials refers to the process of identifying and enrolling participants whose health profiles align with the specific eligibility criteria of a given clinical study. Efficient patient matching is crucial to accelerating the drug development process and reducing the overall cost of clinical trials. It provides an opportunity for participants to access experimental treatments that could potentially improve their health outcomes and, in some cases, transform their quality of life.
Patient matching remains a challenging task despite its critical role in clinical trials. Traditional automated methods, which are mostly rule-based, depend heavily on hand-crafted rules and are often time-consuming (Brøgger-Mikkelsen et al., 2020). These approaches are also highly sensitive to variations in how patient and trial data are expressed in natural language across different sources. Simple machine learning models often struggle with the load of information and the low density of relevant sentences in lengthy electronic health record(EHR) data. They usually underperform when dealing with complex eligibility criteria or inconsistent data formats, resulting in inefficiencies and mismatches that can negatively impact trial outcomes (Kadam et al., 2016).
The emergence of of Large Language Models (LLMs) offers a promising opportunity to improve patient matching. LLMs, such as GPT-4 Achiam et al. (2023), usually possess a large text window and advanced natural language understanding capabilities (Karanikolas et al., 2023), which enable them process complex trail criteria and patient data effectively. However, their general knowledge may not always be sufficient for knowledge-intensive tasks like patient matching in clinical trials, where domain-specific expertise is required. The knowledge gap inherent in LLMs and misssing information in the pre-defined criteria can limit their performance, highlighting the need for LLM augmentation of specialized medical knowledge (Shuster et al., 2021). While research indicates that LLMs do encode medical knowledge in a parameterized form (Singhal et al., 2023), this information can be outdated or misaligned with the special needs of certain clinical trials.
To address these limitations, we present a new framework called: Multi-Agents for Knowledge Augmentation (MAKA). MAKA is a novel architecture designed to enhance the patient matching process by dynamically integrating domain-specific knowledge aligned with the unique context of clinical trials. It consists of five main components: the Knowledge Probing Agent, Navigation Agent, Knowledge Augmentation Agent, Supervision Agent and Matching Agent. By leveraging this multi-agent framework, MAKA effectively identifies and fills knowledge gaps, enhancing the LLM’s ability to understand complex eligibility criteria and patient profiles. This leads to an average improvement of 3% on accuracy and 6% on F1-score across various patient selection criteria, with up to a 10% boost in more knowledge-intensive cases. Overall, our framework shows significant potentials for achieving more accurate patient matching.
2 Related Work
2.1 Clinical Trail Patient Matching
Traditional patient matching systems rely on query-based structured databases, such as Criteria2Query Yuan et al. (2019), or handcrafted rule-based approaches like RBC Oleynik et al. (2019). However, these methods struggle to handle unstructured free-text data and are highly sensitive to variations in formatting and verbal expressions. Unfortunately, a significant portion of EHRs consists of free-text, including clinical notes, surgical records, radiology reports, pathology reports and so on Sun et al. (2018). Thus, the effectiveness and scalability of traditional methods are significantly limited, making them unsuitable for future patient matching tasks.
Recent advancements in LLMs have shown remarkable capabilities in directly processing free-text clinical trial criteria, offering scalable solutions without the need for extensive rule-based preprocessing. Many Studies have applied these models to patient-centric trial matching tasks using real-world data. For instance, Nievas et al. (2024) fine-tuned the open-source Llama 2 Touvron et al. (2023) to automate patient matching. However, fine-tuning LLMs is computationally expensive, prompting alternative approaches that leverage frozen models. Jin et al. (2023) introduced TrialGPT, a patient-centric matching framework that assesses patient eligibility on a criterion-by-criterion basis, filtering out ineligible trials and ranking potentially eligible ones. Their results indicated a strong correlation between TrialGPT’s predictions and expert clinician judgments, achieving accuracy at the criterion level similar to human experts. Similarly, Gupta et al. (2024) proposed OncoLLM, an end-to-end pipeline that converts trial criteria into conversational prompts and uses a Chain-of-Thought (CoT) Wei et al. (2022) reasoning strategy. OncoLLM outperforms GPT-3.5 while underperforms GPT-4, despite its smaller model size.
In contrast, other studies have focused on trial-centric patient matching, typically using public deidentified datasets due to HIPAA-related data constraints. Wornow et al. (2024) developed a LLM-based system with a two-stage retrieval pipeline to identify suitable patients for trials using a zero-shot prompt approach. It reduces token processing costs while outperforming existing state-of-the-art models on the n2c2 2018 cohort selection benchmark Stubbs et al. (2019). Likewise, Beattie et al. (2024) applied a CoT strategy in a Retrieval-Augmented Generation (RAG) Lewis et al. (2020) system using GPT-3.5 Turbo and GPT-4, demonstrating results comparable to those of Wornow et al. (2024).
These works collectively highlight the potential of LLMs in clinical trial matching. However, existing approaches directly employ LLMs without verifying whether the trial criteria are sufficiently understood by the model Webson and Pavlick (2021) or whether there are gaps in the model’s knowledge Lewis et al. (2020). Our work fits within the trial-centric patient matching literature and makes a novel contribution by automatically identifying and addressing these gaps, thereby enhancing the accuracy patient matching.
2.2 LLM-based Multi-Agents (LLM-MA)
LLM-MA have seen rapid development across various domains, driven by the promising potential of LLMs as autonomous agents. Guo et al. (2024) provide a comprehensive overview of recent advancements in LLM-MA applications, categorizing them into two primary areas: Problem-Solving and World Simulation. Problem-solving LLM-MA systems explore the capabilities of agents with specialized expertise, enabling them to outperform single-agent LLM systems through collaborative efforts. For instance, Wang et al. (2024) proposed Mixture-of-Agents, a layered multi-agent architecture that achieves state-of-the-art performance in benchmarks like AlpacaEval 2.0 Dubois et al. (2024). Similarly, Hong et al. (2023) introduced MetaGPT, a meta-programming framework that integrates human-like task decomposition and workflows, outperforming GPT-4 in various code generation tasks.
In the healthcare domain, the use of LLM-MA systems is emerging. TRIAGEAGENT Lu et al. (2024) utilizes LLMs for role-playing interactions, incorporating self-confidence and early-stopping mechanisms in multi-round discussions to improve document reasoning and classification precision for triage tasks. Tang et al. (2023) uses multiple LLM-based agents as domain experts to collaboratively analyze medical reports and reach consensus on medical diagnoses. To the best of our knowledge, this is the first study to leverage LLM-MA for cohort selection in clinical trials. Our framework leverages LLM-MA agency to dynamically augment patient-trial matching with tailored knowledge, thereby enriching the contextual understanding of patient profiles and achieving better patient matching results.
3 Methods
We automate decomposed patient-matching workflows by utilizing the ability of LLMs to process and generate complex natural language. The proposed framework, illustrated in Figure 1, comprises five interconnected components, each with a distinct role in ensuring robust automation.
Knowledge Probing Agent: This component identifies knowledge gaps, both in terms of missing information in criteria and deficiencies in the LLM’s understanding of clinical trials and patient characteristics. By employing prompt-based self-probing techniques, inspired by SELFREF Kadavath et al. (2022) and MOREINFO Feng et al. (2023), the agent evaluates its own responses for correctness and consistency.If the agent determines that the original criterion is detailed and understandable enough for other agents, the criterion will be directly forwarded to the Matching Agent. Otherwise, the criterion is forwarded to the Navigation Agent for further processing.
Navigation Agent: In the MAKA framework, multiple specialized augmentation agents are available to enrich the criterion with domain-specific knowledge. The Navigation Agent determines the most suitable augmentation agent for the task and routes the workflow accordingly.
Knowledge Augmentation Agent: This component enriches the criterion with relevant, domain-specific information from different source. For instance, the Retrieval Agent Lewis et al. (2020) retrieves information from indexed medical databases, the Online Search Agent gathers supplementary knowledge from online sources, and the Self-Augment Agent leverages the LLM’s intrinsic knowledge to refine the original criterion.
Supervision Agent: Acting as a supervisor, this agent ensures that the augmented criterion aligns with the original instructions. For example, the augmented output must retain verbatim elements from the original criterion. This supervisory mechanism enhances the framework’s instruction-following capability and maintains consistency.
Matching Agent: Finally, with either the augmented or unaugmented criterion, as determined by the Knowledge Probing Agent, the Matching Agent evaluates patient eligibility in a zero-shot manner.
4 Evaluation
| Criteria | Accuracy | Precision | Recall | F1-score | ||||||||
| Zero-shot | CoT | MAKA | Zero-shot | CoT | MAKA | Zero-shot | CoT | MAKA | Zero-shot | CoT | MAKA | |
| ABDOMINAL | 0.858 | 0.868 | 0.847 | 0.893 | 0.897 | 0.944 | 0.701 | 0.729 | 0.626 | 0.785 | 0.804 | 0.753 |
| ADVANCED-CAD | 0.854 | 0.854 | 0.868 | 0.895 | 0.872 | 0.929 | 0.853 | 0.882 | 0.841 | 0.873 | 0.877 | 0.883 |
| ALCOHOL-ABUSE | 0.969 | 0.969 | 0.976 | 0.529 | 0.545 | 0.667 | 0.900 | 0.600 | 0.600 | 0.667 | 0.571 | 0.632 |
| ASP-FOR-MI | 0.861 | 0.875 | 0.885 | 0.899 | 0.908 | 0.919 | 0.930 | 0.939 | 0.939 | 0.915 | 0.923 | 0.929 |
| CREATININE | 0.802 | 0.802 | 0.885 | 0.664 | 0.667 | 0.812 | 0.934 | 0.925 | 0.896 | 0.776 | 0.775 | 0.852 |
| DIETSUPP-2MOS | 0.764 | 0.781 | 0.785 | 0.704 | 0.717 | 0.721 | 0.940 | 0.953 | 0.953 | 0.805 | 0.818 | 0.821 |
| DRUG-ABUSE | 0.941 | 0.920 | 0.993 | 0.469 | 0.395 | 0.882 | 1.000 | 1.000 | 1.000 | 0.638 | 0.566 | 0.938 |
| ENGLISH | 0.976 | 0.986 | 0.983 | 0.985 | 0.989 | 0.996 | 0.989 | 0.996 | 0.985 | 0.987 | 0.992 | 0.991 |
| HBA1C | 0.840 | 0.830 | 0.878 | 0.722 | 0.702 | 0.807 | 0.892 | 0.902 | 0.863 | 0.798 | 0.790 | 0.834 |
| KETO-1YR | 0.993 | 0.990 | 0.993 | - | - | - | - | - | - | - | - | - |
| MAJOR-DIABETES | 0.833 | 0.826 | 0.847 | 0.851 | 0.831 | 0.873 | 0.840 | 0.853 | 0.840 | 0.845 | 0.842 | 0.856 |
| MAKES-DECISIONS | 0.920 | 0.924 | 0.931 | 0.985 | 0.985 | 0.996 | 0.931 | 0.935 | 0.931 | 0.957 | 0.959 | 0.963 |
| MI-6MOS | 0.885 | 0.872 | 0.927 | 0.434 | 0.396 | 0.571 | 0.885 | 0.808 | 0.769 | 0.582 | 0.532 | 0.656 |
| Average | 0.884 | 0.883 | 0.909 | 0.727 | 0.715 | 0.822 | 0.894 | 0.870 | 0.846 | 0.785 | 0.769 | 0.828 |
4.1 Data Description
We use the dataset from Track 1 of the 2018 n2c2 challenge Stubbs et al. (2019), which is commonly used for evaluating patient matching in clinical trials. The dataset consists of 288 de-identified patients, each having between 2 to 5 unstructured longitudinal clinical notes in American English. The entire corpus contains 781,006 tokens, with an average of 2,711 tokens per patient. The challenge simulates a synthetic clinical trial with 13 predefined inclusion criteria, resulting in 3744 patient-criteria pairs in total. The detailed criteria can be found in Table 4. For each patient-criteria pair, a binary label is provided, indicating whether the patient "MET" or "NOT MET" the criterion. Two annotators with medical background labeled these data. The disagreements between the two annotators were resolved by an independent MD to establish the final ground truth labels.
4.2 Baseline and Metrics
Previous research has demonstrated the superiority of LLM-based methods over traditional rule-based systems; hence, this study focuses exclusively on LLM-based benchmarks. We include two approaches: the zero-shot method with hand-crafted criteria Wornow et al. (2024), as shown in Table 5, and a Chain-of-Thought (CoT) approach Beattie et al. (2024) using the original criteria. In line with prior literature, we report accuracy, precision, recall, and F1-score for this binary classification task.
Previous studies have primarily focused on evaluating classification performance at the individual criterion level, often overlooking the overall patient-trial assessment. In this study, we address this gap by examining patient-trial selection performance. Specifically, a patient is eligible for the trial only if all inclusion criteria are met. However, no patient in the n2c2 dataset satisfied all 13 criteria. To mitigate this, we synthesize a trail with selected criteria that had at least 100 patients meeting them, resulting in 28 eligible patients from a total cohort of 288. We then evaluated our method’s on this new synthetic trail in identifying eligible patients and reported the findings.
4.3 Results
4.3.1 Criterion-Level Performance
Table 1 shows the performance of three approaches—Zero-shot,CoT, MAKA—across multiple individual criteria. Our results indicate that while the CoT method slightly outperforms Zero-shot in Accuracy and Recall, MAKA consistently achieves higher scores, particularly in Accuracy (0.909) and Precision(0.822) on average. The F1-score of MAKA (0.828) suggests a balanced performance across criteria, although it falls slightly behind in a few specific cases such as "ABDOMINAL" and "English".
The MAKA-augmented criteria, presented in Table 6, demonstrate significant improvements over the baseline human-redefined criteria outlined in Table 5 from Wornow et al. (2024). These enhancements primarily focus on providing explicit clarifications and detailed expansions of key concepts, enabling LLMs to better align with trial requirements. For example, MAKA supplements normal ranges for criteria like "HBA1C" and "CREATININE" across diverse demographic groups. And they are also the criteria that benefit the most from MAKA’s augmentation. Additionally, MAKA provides a more comprehensive definition for criteria like "MAJOR-DIABETES," and use more widely recognized terminology, "heart attack", to explain clinical terms "Myocardial Infarction (MI)". These refinements enhance the ability of LLMs to match patients from clinical notes written in a variety of styles, resulting in notable performance gains for these criteria.
| Metric | Zero-shot | CoT | MAKA |
| Accuracy | 0.8924 | 0.9201 | 0.9306 |
| Precision | 0.4615 | 0.5676 | 0.6333 |
| Recall | 0.6429 | 0.7500 | 0.6786 |
| F1-Score | 0.5373 | 0.6462 | 0.6552 |
4.3.2 Trial-Level Performance
The result of trail-levle analysis can be found in Table 2. At the trial level, where a patient’s eligibility is determined by satisfying all selected criteria, MAKA outperforms both Zero-shot and CoT. MAKA achieves the highest Accuracy (0.9306), Precision (0.6333)and F1-score (0.6552) but underperform CoT on Recall. The trail-level results further validate MAKA’s capability in maintaining a higher balance between identifying eligible patients and minimizing patient-filtering load
Our findings suggest that MAKA effectively can effectively identifies and addresses knowledge gaps in LLMs for trial-centric patient matching. By leveraging agent-augmented criteria, MAKA enhances LLMs’ understanding of clinical trial requirements. And it achieves a better performance on both minimizing manual selection efforts and accurately identifying eligible patients.
5 Limitations
One significant limitation of this study lies in the dataset used for evaluation. The n2c2 dataset, the most commonly used, publicly available dataset for the cohort selection task, is relatively small in size and contains criteria that are neither particularly complex nor challenging for LLM-based processing. As a result, while MAKA is designed as a general framework capable of handling both simple and complex criteria, the dataset’s limitations restrict our ability to thoroughly evaluate its generalizability across a broader range of clinical trial requirements.
An additional observation from our experiments is that the Knowledge Probing Agent predominantly identified gaps in the criteria themselves—primarily issues of clarity or insufficient detail—rather than uncovering knowledge gaps in LLM. Moreover, the Navigation Agent consistently determined that all criteria could be sufficiently augmented by the Self-Augmentation Agent, with no need to employ external retrieval or online search augmentation. These findings suggest that the dataset’s simplicity may have limited the full utilization of MAKA’s modular capabilities.
6 Conclusion
In this paper, we introduced MAKA, a framework designed to automate decomposed patient-matching workflows by leveraging the advanced natural language understanding capabilities of LLMs. The framework’s modular design incorporates five key components—Knowledge Probing Agent, Navigation Agent, Knowledge Augmentation Agent, Supervision Agent, and Matching Agent—each contributing to systematic knowledge augmentation and robust decision-making in patient-trail matching settings.
We demonstrated that MAKA significantly improves patient-matching performance at both the criterion and trial levels. Compared to baseline approaches like Zero-shot and Chain of Thought (CoT), MAKA achieves notable gains in Accuracy, Precision, and F1-score, particularly when applied to complex or ambiguous criteria. By enriching clinical criteria with domain-specific insights, MAKA consistently outperforms alternatives while maintaining a balance between precision and recall. Further investigation of agent behaviors indicates that the performance improvements primarily stem from filling the knowledge gaps existing in the criteria, rather than in LLMs.
Our results highlight the practical benefits of augmenting LLM capabilities with a structured muti-agent workflow for knowledge enhancement. For example, MAKA-augmented criteria improves understanding of clinical trial requirements by aligning better with diverse patient note styles and heterogeneity of criterion. The results validate MAKA’s potential for real-world deployment in clinical trial workflows.
It is worth noting that the design of MAKA supports the future development of a human-AI collaboration system. For the trial-centric patient matching task, the volume of criterion-supervision performed by the supervisory agent is relatively small and can be efficiently handled by human experts with minimal time investment. Exploring how humans can further enhance and facilitate this system presents an interesting direction for future research.
To address the limitations of this study and validate MAKA’s generalizability, future work could also focus on evaluating the framework using more challenging datasets, such as those involving oncology trial matching which contains highly complex criteria or those with a cohort of larger size. Testing MAKA in thess contexts would not only provide a more rigorous assessment of its capabilities but also demonstrate its potential to address critical real-world challenges in clinical trial recruitment.
Acknowledgments
We gratefully acknowledge support from the OpenAI Research Access Program, which provided access to advanced language models that were instrumental in developing and evaluating the framework presented in this work. This support significantly facilitated the experiments and analyses conducted for this research.
Appendix A Criteria
A.1 N2C2 Criteria Statistics
| Criterion | Met Count | Not Met Count |
| ABDOMINAL | 107 | 181 |
| ADVANCED-CAD | 170 | 118 |
| ALCOHOL-ABUSE | 10 | 278 |
| ASP-FOR-MI | 230 | 58 |
| CREATININE | 106 | 182 |
| DIETSUPP-2MOS | 149 | 139 |
| DRUG-ABUSE | 15 | 273 |
| ENGLISH | 265 | 23 |
| HBA1C | 102 | 186 |
| KETO-1YR | 1 | 287 |
| MAJOR-DIABETES | 156 | 132 |
| MAKES-DECISIONS | 277 | 11 |
| MI-6MOS | 26 | 262 |
A.2 N2C2 Original Criteria
| Criteria | Definition |
| DRUG-ABUSE | Drug abuse, current or past. |
| ALCOHOL-ABUSE | Current alcohol use over weekly recommended limits. |
| ENGLISH | Patient must speak English. |
| MAKES-DECISIONS | Patient must make their own medical decisions. |
| ABDOMINAL | History of intra-abdominal surgery, small or large intestine resection, or small bowel obstruction. |
| MAJOR-DIABETES | Major diabetes-related complications, such as: amputation, kidney damage, skin conditions, retinopathy, nephropathy, neuropathy. |
| ADVANCED-CAD | Advanced cardiovascular disease, defined as two or more of the following: medications for CAD, history of myocardial infarction, angina, ischemia. |
| MI-6MOS | Myocardial infarction in the past 6 months. |
| KETO-1YR | Diagnosis of ketoacidosis in the past year. |
| DIETSUPP-2MOS | Taken a dietary supplement (excluding Vitamin D) in the past 2 months. |
| ASP-FOR-MI | Use of aspirin to prevent myocardial infarction. |
| HBA1C | Any HbA1c value between 6.5 and 9.5%. |
| CREATININE | Serum creatinine > upper limit of normal. |
A.3 Redefined Criteria
| Criteria | Definition |
| ABDOMINAL | History of intra-abdominal surgery. This could include any form of intra-abdominal surgery, including but not limited to small/large intestine resection or small bowel obstruction. |
| ADVANCED-CAD | Advanced cardiovascular disease (CAD). Defined as having 2 or more of the following: (a) Taking 2 or more medications to treat CAD (b) History of myocardial infarction (MI) (c) Currently experiencing angina (d) Ischemia, past or present. The patient must have at least 2 of these categories (a,b,c,d) to meet this criterion, otherwise the patient does not meet this criterion. For ADVANCED-CAD, be strict in your evaluation of the patient – if they just have cardiovascular disease, then they do not meet this criterion. |
| ALCOHOL-ABUSE | Current alcohol use over weekly recommended limits. |
| ASP-FOR-MI | Use of aspirin for preventing myocardial infarction (MI). |
| CREATININE | Serum creatinine level above the upper normal limit. |
| DIETSUPP-2MOS | Consumption of a dietary supplement (excluding vitamin D) in the past 2 months. To assess this criterion, go through the list of medications and supplements taken from the note. If a substance could potentially be used as a dietary supplement (i.e., it is commonly used as a dietary supplement, even if it is not explicitly stated as being used as a dietary supplement), then the patient meets this criterion. Be lenient and broad in what is considered a dietary supplement. For example, a ’multivitamin’ and ’calcium carbonate’ should always be considered a dietary supplement if they are included in this list. |
| DRUG-ABUSE | Current or past history of drug abuse. |
| ENGLISH | Patient speaks English. Assume that the patient speaks English, unless otherwise explicitly noted. If the patient’s language is not mentioned in the note, then assume they speak English and thus meet this criteria. |
| HBA1C | Any hemoglobin A1c (HbA1c) value between 6.5% and 9.5%. |
| KETO-1YR | Diagnosis of ketoacidosis within the past year. |
| MAJOR-DIABETES | Major diabetes-related complication. Examples of “major complication” (as opposed to “minor complication”) include, but are not limited to, any of the following that are a result of (or strongly correlated with) uncontrolled diabetes: • Amputation • Kidney damage • Skin conditions • Retinopathy • Nephropathy • Neuropathy. Additionally, if multiple conditions together imply a severe case of diabetes, then count that as a major complication. |
| MAKES-DECISIONS | Patient must make their own medical decisions. Assume that the patient makes their own medical decisions, unless otherwise explicitly noted. There is no information provided about the patient’s ability to make their own medical decisions, then assume they do make their own decisions and therefore meet this criteria. |
| MI-6MOS | Myocardial infarction (MI) within the past 6 months. |
A.4 MAKA-Augemented Criteria
| Criteria | Definition |
| ABDOMINAL | Criteria: History of intra-abdominal surgery, including small/large intestine resection or small bowel obstruction. |
| Explanation: Intra-abdominal surgery involves operations on organs within the abdominal cavity (e.g., stomach, intestines). Small bowel obstruction refers to blockages in the small intestine that may require surgery. | |
| ADVANCED-CAD | Criteria: Advanced cardiovascular disease (CAD) defined by having two or more of the following: (a) Taking 2 or more medications for CAD, (b) History of myocardial infarction (MI), (c) Current angina, or (d) Ischemia (past or present). |
| Explanation: CAD affects heart function due to blocked arteries. Definitions include: MI (heart attack), angina (chest pain), ischemia (reduced blood flow), and specific medications like statins or beta-blockers. A diagnosis of CAD alone does not qualify as advanced. | |
| ALCOHOL-ABUSE | Criteria: Current alcohol use exceeding weekly recommended limits. |
| Explanation: Weekly limits: up to 14 standard drinks. A standard drink is 14g of pure alcohol (12oz beer, 5oz wine, or 1.5oz spirits). | |
| ASP-FOR-MI | Criteria: Use of aspirin to prevent myocardial infarction (MI). |
| Explanation: Aspirin reduces MI risk by preventing clots but requires medical advice to balance benefits (risk reduction) and potential side effects (e.g., bleeding). | |
| CREATININE | Criteria: Serum creatinine level above the upper normal limit. |
| Explanation: Creatinine reflects kidney function. Normal levels vary by age and gender (e.g., 0.6-1.2 mg/dL for adult males). Elevated levels suggest potential kidney impairment. | |
| DIETSUPP-2MOS | Criteria: Consumption of a dietary supplement (excluding Vitamin D) in the past 2 months. |
| Explanation: Supplements include multivitamins, calcium carbonate, fish oil, or herbal supplements like ginseng. Excludes Vitamin D explicitly. | |
| DRUG-ABUSE | Criteria: Current or past drug abuse history. |
| Explanation: Drug abuse includes misuse of psychoactive substances, with potential consequences like health issues or legal problems. | |
| ENGLISH | Criteria: Patient must speak English. |
| Explanation: English proficiency ensures effective participation in English-based trials. Assume English proficiency unless noted otherwise. | |
| HBA1C | Criteria: Hemoglobin A1c (HbA1c) between 6.5% and 9.5%. |
| Explanation: HbA1c measures average blood sugar over 2–3 months. The range identifies diabetes patients with moderate control, excluding poorly controlled cases. | |
| KETO-1YR | Criteria: Diagnosis of ketoacidosis in the past year. |
| Explanation: Ketoacidosis is a serious condition with high blood ketones, often related to uncontrolled diabetes. Symptoms include nausea, confusion, and rapid breathing. | |
| MAJOR-DIABETES | Criteria: Major diabetes-related complications like amputation, kidney damage, or neuropathy. |
| Explanation: Includes significant health effects (e.g., nephropathy, retinopathy) stemming from uncontrolled diabetes. Multiple minor complications combined may also qualify. | |
| MAKES-DECISIONS | Criteria: Patient must make their own medical decisions. |
| Explanation: This assesses cognitive ability to understand medical options, communicate choices, and appreciate consequences. | |
| MI-6MOS | Criteria: Myocardial infarction (MI) within the past 6 months. |
| Explanation: MI refers to heart muscle damage from blocked blood flow. Assess via medical history and records confirming the event within 6 months. |

### v2

@@
-Enhancing Clinical Trial Patient Matching through Knowledge Augmentation with Multi-Agents
-Abstract
+Enhancing Clinical Trial Patient Matching through Knowledge Augmentation and Reasoning with Multi-Agents
+Abstract.
 Matching patients effectively and efficiently for clinical trials is a significant challenge due to the complexity and variability of patient profiles and trial criteria.
-This paper presents a novel framework, Multi-Agents for Knowledge Augmentation (MAKA), designed to enhance patient-trial matching by dynamically supplementing matching prompts with external, domain-specific knowledge.
-The MAKA architecture consists of five key components: a knowledge probing agent that detects gaps in domain knowledge, a navigation agent that manages interactions among multiple specialized knowledge augmentation agents, a knowledge augmentation agent that incorporates relevant information into patient-trial matching prompts, a supervision agent aligning the outputs from other agents with the instructions and a matching agent making the final selection decision.
-This approach enhances the accuracy and contextual richness of patient matching, addresses inherent knowledge gaps in both trail criteria and large language models (LLMs), and improves the alignment between patient characteristics and the criteria.
-1 Introduction
-Patient matching for clinical trials refers to the process of identifying and enrolling participants whose health profiles align with the specific eligibility criteria of a given clinical study.
+This paper introduces Multi-Agents for Knowledge Augmentation and Reasoning (MAKAR), a novel framework that enhances patient-trial matching by integrating domain-specific knowledge with structured reasoning.
+MAKAR consists of two key modules: the Augmentation Module and the Reasoning Module.
+The Augmentation Module enriches trial criteria by incorporating detailed explanations of relevant concepts to ensure clarity and completeness.
+The Reasoning Module then evaluates each health condition, following a structured, step-wise approach to determine eligibility and make the final matching decision.
+This paper enhances patient-trial matching by leveraging the agency and reasoning capabilities of large language models (LLMs) through automated agent interactions, including collaboration, critique, and navigation.
+Experimental results on a public dataset demonstrate that our framework surpasses existing benchmarks, achieving up to an 8% improvement in accuracy for specific criteria.
+Furthermore, in a real-world offline test, MAKAR achieved a 100% accuracy.
+These findings show MAKAR’s potential as a scalable and robust solution for clinical trial patient matching.
+1. Introduction
+Patient matching for clinical trials refers to the process of identifying and enrolling participants whose health profiles meet the specific eligibility criteria of a given clinical study.
 Efficient patient matching is crucial to accelerating the drug development process and reducing the overall cost of clinical trials.
-It provides an opportunity for participants to access experimental treatments that could potentially improve their health outcomes and, in some cases, transform their quality of life.
+It also provides an opportunity for participants to access experimental treatments that could potentially improve their health outcomes and, in some cases, transform their quality of life.
 Patient matching remains a challenging task despite its critical role in clinical trials.
@@
 They usually underperform when dealing with complex eligibility criteria or inconsistent data formats, resulting in inefficiencies and mismatches that can negatively impact trial outcomes (Kadam et al., 2016).
-The emergence of of Large Language Models (LLMs) offers a promising opportunity to improve patient matching.
-LLMs, such as GPT-4 Achiam et al.
-(2023), usually possess a large text window and advanced natural language understanding capabilities (Karanikolas et al., 2023), which enable them process complex trail criteria and patient data effectively.
+The emergence of of LLMs offers a promising opportunity to improve patient matching.
+LLMs, such as GPT-4 (Achiam et al., 2023), usually possess a large text window and advanced natural language understanding capabilities (Karanikolas et al., 2023), which enable them process complex trial criteria and patient data effectively.
 However, their general knowledge may not always be sufficient for knowledge-intensive tasks like patient matching in clinical trials, where domain-specific expertise is required.
@@
 While research indicates that LLMs do encode medical knowledge in a parameterized form (Singhal et al., 2023), this information can be outdated or misaligned with the special needs of certain clinical trials.
-To address these limitations, we present a new framework called: Multi-Agents for Knowledge Augmentation (MAKA).
-MAKA is a novel architecture designed to enhance the patient matching process by dynamically integrating domain-specific knowledge aligned with the unique context of clinical trials.
-It consists of five main components: the Knowledge Probing Agent, Navigation Agent, Knowledge Augmentation Agent, Supervision Agent and Matching Agent.
-By leveraging this multi-agent framework, MAKA effectively identifies and fills knowledge gaps, enhancing the LLM’s ability to understand complex eligibility criteria and patient profiles.
-This leads to an average improvement of 3% on accuracy and 6% on F1-score across various patient selection criteria, with up to a 10% boost in more knowledge-intensive cases.
-Overall, our framework shows significant potentials for achieving more accurate patient matching.
-2 Related Work
-2.1 Clinical Trail Patient Matching
-Traditional patient matching systems rely on query-based structured databases, such as Criteria2Query Yuan et al.
-(2019), or handcrafted rule-based approaches like RBC Oleynik et al.
-(2019).
+To address these limitations, we present a new framework called: Multi-Agents for Knowledge Augmentation and Reasoning (MAKAR).
+MAKAR is a novel framework designed to enhance the patient matching process.
+The framework consists of two core modules: the Augmentation Module and the Reasoning Module.
+The Augmentation Module enhances trial criteria by providing detailed explanations and includes three key components: a Router Agent, multiple Augmentation Agents, and a Critic Agent.
+The Reasoning Module systematically evaluates each criterion-related condition in a step-by-step manner to determine eligibility and make the final matching decision.
+It comprises a Reasoning Agent and a Supervision Matching Agent.
+By leveraging this multi-agent framework, MAKAR improves the performance on patient-trial matching by dynamically incorporating domain-specific knowledge and step-wise reasoning.
+Our experimental results demonstrate the effectiveness of MAKAR.
+On a public dataset, our framework achieves an average improvement of 3% over benchmark methods, with up to an 8% improvement for complex criteria.
+Furthermore, in an offline real-world data test, MAKAR outperformed the state-of-the-art closed-source reasoning LLMs and achieved 100% accuracy, showing its reliability in practical applications.
+Our key contributions are as follows:
+We propose an automated method to refine and enhance trial criteria, making them more interpretable and suitable for AI-driven patient-trial matching.
+We introduce a multi-agent framework that combines augmentation and reasoning, achieving strong performance across both public and real-world datasets.
+We identify a gap in existing approaches and highlight the importance of explicitly incorporating well-structured, detailed explanations of related concepts to improve LLM-driven clinical reasoning.
+2. Related Work
+2.1. Clinical Trial Patient Matching
+Traditional patient matching systems rely on query-based structured databases, such as Criteria2Query (Yuan et al., 2019), or handcrafted rule-based approaches like RBC (Oleynik et al., 2019).
 However, these methods struggle to handle unstructured free-text data and are highly sensitive to variations in formatting and verbal expressions.
-Unfortunately, a significant portion of EHRs consists of free-text, including clinical notes, surgical records, radiology reports, pathology reports and so on Sun et al.
-(2018).
+Unfortunately, a significant portion of EHRs consists of free-text, including clinical notes, surgical records, radiology reports, pathology reports and so on (Sun et al., 2018).
 Thus, the effectiveness and scalability of traditional methods are significantly limited, making them unsuitable for future patient matching tasks.
@@
 For instance, Nievas et al.
-(2024) fine-tuned the open-source Llama 2 Touvron et al.
-(2023) to automate patient matching.
+(2024) fine-tuned the open-source Llama 2 (Touvron et al., 2023) to automate patient matching.
 However, fine-tuning LLMs is computationally expensive, prompting alternative approaches that leverage frozen models.
@@
 Similarly, Gupta et al.
-(2024) proposed OncoLLM, an end-to-end pipeline that converts trial criteria into conversational prompts and uses a Chain-of-Thought (CoT) Wei et al.
-(2022) reasoning strategy.
+(2024) proposed OncoLLM, an end-to-end pipeline that converts trial criteria into conversational prompts and uses a Chain-of-Thought (CoT) (Wei et al., 2022) reasoning strategy.
 OncoLLM outperforms GPT-3.5 while underperforms GPT-4, despite its smaller model size.
@@
 (2024) developed a LLM-based system with a two-stage retrieval pipeline to identify suitable patients for trials using a zero-shot prompt approach.
-It reduces token processing costs while outperforming existing state-of-the-art models on the n2c2 2018 cohort selection benchmark Stubbs et al.
-(2019).
+It reduces token processing costs while outperforming existing state-of-the-art models on the n2c2 2018 cohort selection benchmark (Stubbs et al., 2019).
 Likewise, Beattie et al.
-(2024) applied a CoT strategy in a Retrieval-Augmented Generation (RAG) Lewis et al.
-(2020) system using GPT-3.5 Turbo and GPT-4, demonstrating results comparable to those of Wornow et al.
+(2024) applied a CoT strategy in a Retrieval-Augmented Generation (RAG) (Lewis et al., 2020) system using GPT-3.5 Turbo and GPT-4, demonstrating results comparable to those of Wornow et al.
 (2024).
 These works collectively highlight the potential of LLMs in clinical trial matching.
-However, existing approaches directly employ LLMs without verifying whether the trial criteria are sufficiently understood by the model Webson and Pavlick (2021) or whether there are gaps in the model’s knowledge Lewis et al.
-(2020).
+However, existing approaches directly employ LLMs without verifying whether the trial criteria are sufficiently understood by the model (Webson and Pavlick, 2021) or whether there are gaps in the model’s knowledge (Lewis et al., 2020).
 Our work fits within the trial-centric patient matching literature and makes a novel contribution by automatically identifying and addressing these gaps, thereby enhancing the accuracy patient matching.
-2.2 LLM-based Multi-Agents (LLM-MA)
+2.2. LLM-based Multi-Agents (LLM-MA)
 LLM-MA have seen rapid development across various domains, driven by the promising potential of LLMs as autonomous agents.
@@
 For instance, Wang et al.
-(2024) proposed Mixture-of-Agents, a layered multi-agent architecture that achieves state-of-the-art performance in benchmarks like AlpacaEval 2.0 Dubois et al.
-(2024).
-Similarly, Hong et al.
-(2023) introduced MetaGPT, a meta-programming framework that integrates human-like task decomposition and workflows, outperforming GPT-4 in various code generation tasks.
-In the healthcare domain, the use of LLM-MA systems is emerging.
-TRIAGEAGENT Lu et al.
-(2024) utilizes LLMs for role-playing interactions, incorporating self-confidence and early-stopping mechanisms in multi-round discussions to improve document reasoning and classification precision for triage tasks.
-Tang et al.
-(2023) uses multiple LLM-based agents as domain experts to collaboratively analyze medical reports and reach consensus on medical diagnoses.
-To the best of our knowledge, this is the first study to leverage LLM-MA for cohort selection in clinical trials.
-Our framework leverages LLM-MA agency to dynamically augment patient-trial matching with tailored knowledge, thereby enriching the contextual understanding of patient profiles and achieving better patient matching results.
-3 Methods
-We automate decomposed patient-matching workflows by utilizing the ability of LLMs to process and generate complex natural language.
-The proposed framework, illustrated in Figure 1, comprises five interconnected components, each with a distinct role in ensuring robust automation.
-Knowledge Probing Agent: This component identifies knowledge gaps, both in terms of missing information in criteria and deficiencies in the LLM’s understanding of clinical trials and patient characteristics.
-By employing prompt-based self-probing techniques, inspired by SELFREF Kadavath et al.
-(2022) and MOREINFO Feng et al.
-(2023), the agent evaluates its own responses for correctness and consistency.If the agent determines that the original criterion is detailed and understandable enough for other agents, the criterion will be directly forwarded to the Matching Agent.
-Otherwise, the criterion is forwarded to the Navigation Agent for further processing.
-Navigation Agent: In the MAKA framework, multiple specialized augmentation agents are available to enrich the criterion with domain-specific knowledge.
-The Navigation Agent determines the most suitable augmentation agent for the task and routes the workflow accordingly.
-Knowledge Augmentation Agent: This component enriches the criterion with relevant, domain-specific information from different source.
-For instance, the Retrieval Agent Lewis et al.
-(2020) retrieves information from indexed medical databases, the Online Search Agent gathers supplementary knowledge from online sources, and the Self-Augment Agent leverages the LLM’s intrinsic knowledge to refine the original criterion.
-Supervision Agent: Acting as a supervisor, this agent ensures that the augmented criterion aligns with the original instructions.
-For example, the augmented output must retain verbatim elements from the original criterion.
-This supervisory mechanism enhances the framework’s instruction-following capability and maintains consistency.
-Matching Agent: Finally, with either the augmented or unaugmented criterion, as determined by the Knowledge Probing Agent, the Matching Agent evaluates patient eligibility in a zero-shot manner.
-4 Evaluation
+(2024) proposed Mixture-of-Agents, a layered multi-agent architecture that achieves state-of-the-art performance in benchmarks like AlpacaEval 2.0 (Dubois et al., 2024).
+In the healthcare domain, the application of LLM-MA systems is rapidly gaining traction.
+Yue et al.(Yue and Fu, 2024) introduce ClinicalAgent, a multi-agent architecture powered by LLM reasoning for multi-agent collaboration, achieving higher precision in predicting clinical trial outcomes.
+Similarly, Liao et al.(Liao et al., 2024) propose REFLECTool, a multi-agent framework that enables agents to learn how to select domain-specific tools, thereby extend their capabilities.
+TriageAgent(Lu et al., 2024) leverages LLMs for role-playing interactions, integrating self-confidence assessments and early-stopping mechanisms in multi-round discussions to improve document reasoning and classification accuracy for triage tasks.
+Tang et al.(Tang et al., 2023) use multiple LLM-based agents as domain experts to collaboratively analyze medical reports.
+And Li et al.
+(Li et al., 2024) further introduce the multi-agent collaboration framework to handle multimodal tasks.
+To the best of our knowledge, this study is the first to apply an LLM-MA framework for cohort selection in clinical trials.
+Our approach explore the agency and reasoning capabilities of LLM-MA to dynamically enhance patient-trial matching with tailored knowledge.
+Our approach allows a better contextual understanding of patient profiles and lead to improved matching performance and more informed decision-making in clinical trial recruitment.
 | Criteria | Accuracy | Precision | Recall | F1-score | ||||||||
-| Zero-shot | CoT | MAKA | Zero-shot | CoT | MAKA | Zero-shot | CoT | MAKA | Zero-shot | CoT | MAKA | |
+| Zero-shot | CoT | MAKAR | Zero-shot | CoT | MAKAR | Zero-shot | CoT | MAKAR | Zero-shot | CoT | MAKAR | |
 | ABDOMINAL | 0.858 | 0.868 | 0.847 | 0.893 | 0.897 | 0.944 | 0.701 | 0.729 | 0.626 | 0.785 | 0.804 | 0.753 |
@@
 | MI-6MOS | 0.885 | 0.872 | 0.927 | 0.434 | 0.396 | 0.571 | 0.885 | 0.808 | 0.769 | 0.582 | 0.532 | 0.656 |
-| Average | 0.884 | 0.883 | 0.909 | 0.727 | 0.715 | 0.822 | 0.894 | 0.870 | 0.846 | 0.785 | 0.769 | 0.828 |
-4.1 Data Description
-We use the dataset from Track 1 of the 2018 n2c2 challenge Stubbs et al.
-(2019), which is commonly used for evaluating patient matching in clinical trials.
+| Average | 0.884 | 0.884 | 0.908 | 0.750 | 0.740 | 0.841 | 0.899 | 0.876 | 0.853 | 0.801 | 0.786 | 0.841 |
+3. Methods
+The proposed framework, shown in Figure 1, consists of two core modules: the Augmentation Module and the Reasoning Module.
+The augmentation module is designed to generate criteria with comprehensive explanations of all relevant concepts, while the reasoning module evaluates each condition within the criteria to determine eligibility and make the final matching decision.
+Additionally, the framework contains five interconnected components, each outlined below.
+Router Agent In the MAKAR framework, multiple specialized augmentation agents are available to augment the criterion with domain-specific knowledge.
+The Router Agent determines the most suitable augmentation agent for the task and routes the workflow accordingly.
+Inspired by SELFREF (Kadavath et al., 2022) and MOREINFO (Feng et al., 2023), we adopt a prompt-based self-probing strategy.
+If the agent determines that the original criterion is detailed and understandable enough for other agents, the criterion will be directly forwarded to the Reasoning Module.
+Otherwise, the criterion is forwarded to the Knowledge Augmentation Agent chosen by the Router Agent for further processing.
+Knowledge Augmentation Agent: This component augments the criterion with relevant, domain-specific information from different source.
+For instance, the Retrieval Agent (Lewis et al., 2020) retrieves information from indexed medical databases, the Online Search Agent gathers supplementary knowledge from online sources, and the Self-Augment Agent leverages the LLM’s intrinsic knowledge to refine the original criterion.
+Critic Agent: Acting as a supervisor, this agent ensures that the augmented criterion is aligned with the original instructions.
+For instance, the original criterion must be preserved verbatim.
+Additionally, the critic agent verifies that the augmented output is well-structured (e.g., concepts are clearly itemized) to facilitate further step-wise reasoning.
+This critic mechanism are designed to improve the framework’s ability to follow instructions accurately while maintaining consistency and clarity.
+Reasoning Agent With the augmented criterion and patient EHR data, the Reasoning Agent evaluates patient eligibility by assessing for each concepts mentioned in the augmented criterion.
+It generates a detailed reasoning summary, including assessments for all relevant health conditions, and forwards it to the Supervision Matching Agent.
+Matching Agent: The Matching Agent determines patient eligibility by reviewing the reasoning assessment alongside the patient’s EHR data, making a final judgment in a zero-shot manner.
+4. Evaluation & Results
+Due to the sensitive nature of patient data, which is strictly protected under The Health Insurance Portability and Accountability Act (HIPAA), and the compliance requirements for processing detail clinical trial documents, full-scale online deployment is not currently feasible—particularly due to concerns about data security when using closed-source LLMs.
+Instead, to evaluate the effectiveness of our methods, we conducted two experiments: (1) an assessment using a public dataset and (2) a limited-scale offline evaluation with deidentified real-world data.
+4.1. Data
+The public dataset we use is from Track 1 of the 2018 n2c2 challenge (Stubbs et al., 2019), which is commonly used for evaluating patient matching in clinical trials.
 The dataset consists of 288 de-identified patients, each having between 2 to 5 unstructured longitudinal clinical notes in American English.
@@
 The detailed criteria can be found in Table 6.
-For each patient-criteria pair, a binary label is provided, indicating whether the patient "MET" or "NOT MET" the criterion.
+For each patient-criteria pair, a binary label is provided, indicating whether the patient ”MET” or ”NOT MET” the criterion.
 Two annotators with medical background labeled these data.
 The disagreements between the two annotators were resolved by an independent MD to establish the final ground truth labels.
-4.2 Baseline and Metrics
+The criteria in the n2c2 dataset are relatively straightforward and the patient records are well-structured, making them easy to follow.
+To evaluate the potential of our proposed method in a more complex scenario, we constructed a real-world dataset.
+Specifically, we sampled 30 eligibility criteria—15 inclusion and 15 exclusion—from ongoing breast cancer trials listed on ClinicalTrials.gov.
+We then manually selected 10 patients who have been diagnosed with breast cancer from our in-house EHR system.
+For each patient, we extracted their most recent deidentified clinical visit record.
+An expert subsequently asseessed whether each patient met each criterion, resulting in 300 labeled criterion-patient pairs.
+Notably, if a patient’s information was insufficient to determine eligibility for a given criterion, they were labeled as ”NOT MET” to ensure a conservative patient selection.
+We listed the sampled criterion in Table 9
+4.2. Baselines and Metrics
 Previous research has demonstrated the superiority of LLM-based methods over traditional rule-based systems; hence, this study focuses exclusively on LLM-based benchmarks.
-We include two approaches: the zero-shot method with hand-crafted criteria Wornow et al.
-(2024), as shown in Table 5, and a Chain-of-Thought (CoT) approach Beattie et al.
-(2024) using the original criteria.
+We include two approaches in Experiment 1: the zero-shot method with hand-crafted criteria (Wornow et al., 2024), as shown in Table LABEL:table:redefined_criteria, and a Chain-of-Thought (CoT) approach (Beattie et al., 2024) using the original criteria.
+In addtion, we also include o3-mini from OpenAI as the benchmark in Experiment 2 since reasoning LLMs have shown significant improvement on complex clinical scenarios(Xie et al., 2024).
 In line with prior literature, we report accuracy, precision, recall, and F1-score for this binary classification task.
-Previous studies have primarily focused on evaluating classification performance at the individual criterion level, often overlooking the overall patient-trial assessment.
-In this study, we address this gap by examining patient-trial selection performance.
+Previous studies using n2c2 dataset have primarily focused on evaluating classification performance at the individual criterion level, often overlooking the overall patient-trial assessment.
+In this study, we also examine patient-trial selection performance in Experiment 1.
 Specifically, a patient is eligible for the trial only if all inclusion criteria are met.
 However, no patient in the n2c2 dataset satisfied all 13 criteria.
-To mitigate this, we synthesize a trail with selected criteria that had at least 100 patients meeting them, resulting in 28 eligible patients from a total cohort of 288.
-We then evaluated our method’s on this new synthetic trail in identifying eligible patients and reported the findings.
-4.3 Results
-4.3.1 Criterion-Level Performance
-Table 1 shows the performance of three approaches—Zero-shot,CoT, MAKA—across multiple individual criteria.
-Our results indicate that while the CoT method slightly outperforms Zero-shot in Accuracy and Recall, MAKA consistently achieves higher scores, particularly in Accuracy (0.909) and Precision(0.822) on average.
-The F1-score of MAKA (0.828) suggests a balanced performance across criteria, although it falls slightly behind in a few specific cases such as "ABDOMINAL" and "English".
-The MAKA-augmented criteria, presented in Table 6, demonstrate significant improvements over the baseline human-redefined criteria outlined in Table 5 from Wornow et al.
-(2024).
-These enhancements primarily focus on providing explicit clarifications and detailed expansions of key concepts, enabling LLMs to better align with trial requirements.
-For example, MAKA supplements normal ranges for criteria like "HBA1C" and "CREATININE" across diverse demographic groups.
-And they are also the criteria that benefit the most from MAKA’s augmentation.
-Additionally, MAKA provides a more comprehensive definition for criteria like "MAJOR-DIABETES," and use more widely recognized terminology, "heart attack", to explain clinical terms "Myocardial Infarction (MI)".
-These refinements enhance the ability of LLMs to match patients from clinical notes written in a variety of styles, resulting in notable performance gains for these criteria.
-| Metric | Zero-shot | CoT | MAKA |
+To mitigate this, we synthesize a trial with selected criteria that had at least 100 patients meeting them, resulting in 28 eligible patients from a total cohort of 288.
+We then evaluated our method’s on this new synthetic trial in identifying eligible patients and reported the findings.
+4.3. Experiment Settings
+Our experiments were conducted on an Amazon EC2 g5.16xlarge instance, equipped with a A10G GPU with 24 GB of memory. Unless otherwise specified, all agents in the framework are powered by GPT-4o.
+For external knowledge augmentation, the Online Search Agent uses the Perplexity AI search engine to fetch up-to-date information from the web.
+The Retrieval Agent uses all-MiniLM-L12-v2, a pre-trained checkpoint from Sentence-Transformer (Reimers, 2019), as its encoder to construct the vector database of MedDRA and FDALabel.
+| Metric | Zero-shot | CoT | MAKAR |
 | Accuracy | 0.8924 | 0.9201 | 0.9306 |
@@
 | F1-Score | 0.5373 | 0.6462 | 0.6552 |
-4.3.2 Trial-Level Performance
-The result of trail-levle analysis can be found in Table 2.
-At the trial level, where a patient’s eligibility is determined by satisfying all selected criteria, MAKA outperforms both Zero-shot and CoT.
-MAKA achieves the highest Accuracy (0.9306), Precision (0.6333)and F1-score (0.6552) but underperform CoT on Recall.
-The trail-level results further validate MAKA’s capability in maintaining a higher balance between identifying eligible patients and minimizing patient-filtering load
-Our findings suggest that MAKA effectively can effectively identifies and addresses knowledge gaps in LLMs for trial-centric patient matching.
-By leveraging agent-augmented criteria, MAKA enhances LLMs’ understanding of clinical trial requirements.
-And it achieves a better performance on both minimizing manual selection efforts and accurately identifying eligible patients.
-5 Limitations
-One significant limitation of this study lies in the dataset used for evaluation.
-The n2c2 dataset, the most commonly used, publicly available dataset for the cohort selection task, is relatively small in size and contains criteria that are neither particularly complex nor challenging for LLM-based processing.
-As a result, while MAKA is designed as a general framework capable of handling both simple and complex criteria, the dataset’s limitations restrict our ability to thoroughly evaluate its generalizability across a broader range of clinical trial requirements.
-An additional observation from our experiments is that the Knowledge Probing Agent predominantly identified gaps in the criteria themselves—primarily issues of clarity or insufficient detail—rather than uncovering knowledge gaps in LLM.
-Moreover, the Navigation Agent consistently determined that all criteria could be sufficiently augmented by the Self-Augmentation Agent, with no need to employ external retrieval or online search augmentation.
-These findings suggest that the dataset’s simplicity may have limited the full utilization of MAKA’s modular capabilities.
-6 Conclusion
-In this paper, we introduced MAKA, a framework designed to automate decomposed patient-matching workflows by leveraging the advanced natural language understanding capabilities of LLMs.
-The framework’s modular design incorporates five key components—Knowledge Probing Agent, Navigation Agent, Knowledge Augmentation Agent, Supervision Agent, and Matching Agent—each contributing to systematic knowledge augmentation and robust decision-making in patient-trail matching settings.
-We demonstrated that MAKA significantly improves patient-matching performance at both the criterion and trial levels.
-Compared to baseline approaches like Zero-shot and Chain of Thought (CoT), MAKA achieves notable gains in Accuracy, Precision, and F1-score, particularly when applied to complex or ambiguous criteria.
-By enriching clinical criteria with domain-specific insights, MAKA consistently outperforms alternatives while maintaining a balance between precision and recall.
-Further investigation of agent behaviors indicates that the performance improvements primarily stem from filling the knowledge gaps existing in the criteria, rather than in LLMs.
-Our results highlight the practical benefits of augmenting LLM capabilities with a structured muti-agent workflow for knowledge enhancement.
-For example, MAKA-augmented criteria improves understanding of clinical trial requirements by aligning better with diverse patient note styles and heterogeneity of criterion.
-The results validate MAKA’s potential for real-world deployment in clinical trial workflows.
-It is worth noting that the design of MAKA supports the future development of a human-AI collaboration system.
-For the trial-centric patient matching task, the volume of criterion-supervision performed by the supervisory agent is relatively small and can be efficiently handled by human experts with minimal time investment.
+4.4. Results of Experiment 1
+4.4.1. Criterion-Level Performance:
+Table 1 shows the performance of three approaches—Zero-shot,CoT, MAKAR—across multiple individual criteria.
+Our results indicate that while the CoT method slightly outperforms Zero-shot in Accuracy and Recall, MAKAR consistently achieves higher scores, particularly in Accuracy (0.909) and Precision(0.822) on average.
+The F1-score of MAKAR (0.828) suggests a balanced performance across criteria, although it falls slightly behind in a few specific cases such as ”ABDOMINAL” and ”English”.
+The MAKAR-augmented criteria, presented in Table LABEL:tab:aug_criteria, demonstrate significant improvements over the baseline human-redefined criteria outlined in Table LABEL:table:redefined_criteria from Wornow et al.
+(2024).
+These enhancements primarily focus on providing explicit clarifications and detailed expansions of key concepts, enabling LLMs to better align with trial requirements.
+For example, MAKAR supplements normal ranges for criteria like ”HBA1C” and ”CREATININE” across diverse demographic groups.
+And they are also the criteria that benefit the most from MAKAR’s augmentation.
+Additionally, MAKAR provides a more comprehensive definition for criteria like ”MAJOR-DIABETES,” and use more widely recognized terminology, ”heart attack”, to explain clinical terms ”Myocardial Infarction (MI)”.
+These refinements improve the ability of LLMs to match patients from clinical notes written in a variety of styles, resulting in notable performance gains for these criteria.
+4.4.2. Trial-Level Performance
+The result of trial-levle analysis can be found in Table 2.
+At the trial level, where a patient’s eligibility is determined by satisfying all selected criteria, MAKAR outperforms both Zero-shot and CoT.
+MAKAR achieves the highest Accuracy (0.9306), Precision (0.6333)and F1-score (0.6552) but underperform CoT on Recall.
+The trial-level results further validate MAKAR’s capability in maintaining a higher balance between identifying eligible patients and minimizing patient-filtering load
+| Inclusion Criteria | |||||
+| Metric | GPT-4o | o1 | o3-mini | CoT | MAKAR |
+| Accuracy | 0.9000 | 0.9333 | 0.8533 | 0.9667 | 1.000 |
+| Precision | 0.7884 | 0.8407 | 0.7211 | 0.9385 | 1.000 |
+| Recall | 0.9020 | 0.9413 | 0.8151 | 0.9208 | 1.000 |
+| F1-Score | 0.8281 | 0.8800 | 0.7516 | 0.9294 | 1.000 |
+| Exclusion Criteria | |||||
+| Metric | GPT-4o | o1 | o3-mini | CoT | MAKARCorr |
+| Accuracy | 0.9000 | 0.9267 | 0.9133 | 0.9467 | 1.000 |
+| Precision | 0.8693 | 0.8987 | 0.8840 | 0.9265 | 1.000 |
+| Recall | 0.9090 | 0.9416 | 0.9253 | 0.9487 | 1.000 |
+| F1-Score | 0.8845 | 0.9153 | 0.8999 | 0.9365 | 1.000 |
+4.5. Results of Experiment 2
+4.5.1. Results
+Table 3 presents the evaluation results of our method on the in-house real-world dataset.
+We compared its performance against zero-shot prompting with GPT-4o, zero-shot prompting with two reasoning models (OpenAI-o3-mini and OpenAI-o1) and CoT strategy with GPT-4o.
+Our method achieves perfect scores across all four metrics.
+Notably, for exclusion criteria, we report results for a corrected version of MAKAR (MAKARCorr).
+In the initial version, the model was not explicitly instructed on whether a given criterion was for inclusion or exclusion.
+Surprisingly, it was able to infer some exclusion criteria but focused on determining a patient’s overall eligibility for a trial rather than assessing whether a specific criterion was met.
+For instance, if a patient was 75 years old, they should have been marked as ”MET” for an exclusion criterion of ”greater than 70 years old”.
+While MAKAR correctly identified the age range and provided a logical reasoning process, it mistakenly output ”Not MET”—interpreting the task as determining overall trial eligibility rather than verifying the criterion itself.
+The corrected version addresses this by explicitly instructing the model to evaluate each criterion independently rather than making a broader eligibility judgment.
+MAKAR outperforms all the benchmarks and the CoT follows closely, achieving 96.67% accuracy and 92.94% F1-score for inclusion criteria and 94.67% accuracy with 93.65% F1-score for exclusion criteria.
+o3-mini, a smaller model, is outperformed by o1 but shows competitive performance on exclusion criteria with zero-shot method with GPT-4o.
+Interestingly, Experiment 2 yields better overall performance than Experiment 1, despite its more knowledge-intensive criteria.
+We hypothesize that this is due to differences in the labeling process—the n2c2 challenge involves more subjective judgment, while our setup explicitly defines insufficient information as ”NOT MET”.
+4.5.2. Error Analysis
+Despite MAKAR achieving 100% accuracy in the second experiment, we manually checked the reasoning results of all 300 patient-criterion pair and identified two reasoning errors that did not affect the final output.
+First, in Exclusion Criterion 13 (Patients with ongoing indications for cardioprotective medication—ACE inhibitors, ARBs, and/or beta-blockers), MAKAR reasoned: ”The patient is on metoprolol, a beta-blocker, suggesting some ongoing indication for its use.
+However, without explicit information on the specific condition necessitating this medication (e.g., arrhythmia, prophylactic use post-cardiac event other than MI, or another reason), the justification remains uncertain.” This reasoning was overly conservative, but the Supervision Matching Agent correctly classified it as MET, aligning with expert judgment.
+Second, in Inclusion Criterion 5 (Histologically confirmed breast cancer with an invasive component 20 mm and/or morphologically confirmed spread to regional lymph nodes(stage cT2-cT4 with any cN, or cN1-cN3 with any cT).), MAKAR reasoned: ”The documentation mentions a suspicious breast nodule at the right breast (1.4 × 1.5 × 0.5 cm), which measures more than 20 mm in one dimension.” This was an obvious error, as the maximum dimension stated is only 1.5 cm.
+The mistake may come from over-reasoning, possibly computing the largest possible diagonal as .
+However, the final decision remained correct since the criterion lymph node involvement still holds.
+4.6. Ablation
+In this section, we validate the contribution of each module in our framework.
+First, we remove the knowledge augmentation module to assess the reasoning module (MAKA-reason) using the original criteria, which essentially mirrors the CoT approach in Table 3.
+Next, we remove the reasoning module and evaluate the knowledge augmentation module (MAKA-aug) independently by performing zero-shot matching with the augmented criteria.
+The results, presented in Table 4, indicate that removing the reasoning module has a greater negative impact on inclusion criteria matching (F1-score: 88.37% vs.
+92.94%), while removing the augmentation module has a larger effect on exclusion criteria (F1-score: 93.65% vs.
+94.41%).
+Although we did not involve any reasoning-LLM-based agents, our approach still shows impressive reasoning capabilities.
+This can be attributed to the well-structured augmented criteria and the CoT strategy employed by the reasoning agent.
+For instance, when evaluating whether a patient meets the Inclusion Criterion 14 (Postmenopausal hormone receptor-positive patients), MAKA reasoned: ”…the record does not explicitly state whether the breast cancer was hormone receptor positive.
+However, the use of tamoxifen, which is specifically prescribed for estrogen receptor-positive breast cancer by blocking estrogen receptors, strongly implies that her breast cancer was hormone receptor positive.” Both MAKA and MAKA-reason successfully made the correct match, while MAKA-aug failed to capture this implicit relationship.
+Additionally, we also observed cases where knowledge augmentation improved performance.
+For example, in Exclusion Criterion 7, it mentions: ”…all histological lesions were HER2 1+ or 2+, not detected by FISH amplification, and HR positive according to ASCO guidelines”.
+MAKA-aug expanded the definitions from American Society of Clinical Oncology(ASCO) guidelines, made correct matches for all 10 patients.
+In contrast, MAKA-reason correctly identified only 8 out of 10.
+| Inclusion Criteria | |||
+| Metric | MAKAR-reason | MAKAR-aug | MAKAR |
+| Accuracy | 0.9667 | 0.9333 | 1.000 |
+| Precision | 0.9385 | 0.8387 | 1.000 |
+| Recall | 0.9208 | 0.9612 | 1.000 |
+| F1-Score | 0.9294 | 0.8837 | 1.000 |
+| Exclusion Criteria | |||
+| Metric | MAKAR-reason | MAKAR-aug | MAKARCorr |
+| Accuracy | 0.9467 | 0.9533 | 1.000 |
+| Precision | 0.9265 | 0.9360 | 1.000 |
+| Recall | 0.9487 | 0.9534 | 1.000 |
+| F1-Score | 0.9365 | 0.9441 | 1.000 |
+5. Discussion
+One interesting observation from our experiments is that the Router Agent consistently determined that all criteria could be augmented by the Self-Augmentation Agent, without requiring external retrieval or online search augmentation.
+This was true even for breast cancer-related criteria, which are typically complex and challenging for individuals without professional training.
+This relates to a limitation of our study, which comes from the dataset used for evaluation.
+Both the n2c2 dataset and the real-world dataset are relatively small and contain criteria that are not particularly challenging for LLM-based processing, allowing all benchmark methods to achieve strong performance.
+While MAKAR is designed as a general framework capable of handling both simple and complex criteria, these dataset constraints limit our ability to explore the knowledge boundaries of LLMs and fully assess the framework’s generalizability across a wider range of clinical trial requirements.
+Moreover, this observation raises a discussion about the specific improvements our framework brings to baseline LLMs.
+While the framework is designed to bridge the knowledge gap between trial criteria and LLMs’ existing knowledge, our experiments indicate that LLMs can already expand and interpret the criteria in our datasets accurately.
+Instead, the primary benefit of knowledge augmentation lies in explicitly organizing relevant concepts and definitions in a well-structured format.
+By clearly outlining these key elements, the augmentation provides a type of breakdown of the subsequent reasoning steps, leading to improved overall performance.
+However, despite the dataset limitations, the effectiveness of the augmentation module remains significant.
+One of its key strengths is its potential to interface with diverse data sources, particularly those that are not publicly available, such as paid subscription-based medical databases or proprietary clinical trial repositories.
+Furthermore, it plays a crucial role in integrating the most up-to-date medical knowledge, including newly developed treatments and evolving FDA guidelines.
+This capability ensures that the framework remains adaptable and relevant, even when dealing with rapidly changing medical landscapes where LLMs alone may lack the latest information.
+By systematically incorporating external knowledge, the augmentation module enhances the contextual relevance of trial matching, making it particularly valuable in real-world applications where timely and comprehensive knowledge retrieval is essential.
+6. Conclusion
+In this study, we introduced MAKAR, a scalable multi-agent framework designed to enhance patient-trial matching by combining knowledge augmentation with structured reasoning.
+By integrating an automated augmentation module with a reasoning module, MAKAR refines complex trial criteria and systematically evaluates patient eligibility.
+Our experimental results demonstrate that MAKAR outperforms existing benchmarks, including the state-of-the-art closed-source reasoning LLMs, and achieve 100% accuracy in a real-world offline test.
+And it also achieves up to an 8% improvement on complex criteria in a public dataset.
+Beyond performance gains, our findings emphasize the importance of explicitly structuring and expanding trial criteria to bridge the knowledge gap in AI-driven clinical reasoning.
+It is worth noting that the design of MAKAR supports the future development of a human-AI collaboration system.
+For the trial-centric patient matching task, the workload of criterion-supervision performed by the critic agent is one-time per trial and relatively small and can be efficiently handled by human experts with minimal time investment.
 Exploring how humans can further enhance and facilitate this system presents an interesting direction for future research.
-To address the limitations of this study and validate MAKA’s generalizability, future work could also focus on evaluating the framework using more challenging datasets, such as those involving oncology trial matching which contains highly complex criteria or those with a cohort of larger size.
-Testing MAKA in thess contexts would not only provide a more rigorous assessment of its capabilities but also demonstrate its potential to address critical real-world challenges in clinical trial recruitment.
-Acknowledgments
+To address the limitations of this study and validate MAKAR’s generalizability, future work could also focus on evaluating the framework using more challenging datasets, such as those involving newly launched oncology trial that required up-to-date medical knowledge.
+Testing MAKAR in thess contexts would not only provide a more rigorous assessment of its capabilities but also demonstrate its potential to address critical real-world challenges in clinical trial recruitment.
+7. Acknowledgments
 We gratefully acknowledge support from the OpenAI Research Access Program, which provided access to advanced language models that were instrumental in developing and evaluating the framework presented in this work.
-This support significantly facilitated the experiments and analyses conducted for this research.
 Appendix A Criteria
-A.1 N2C2 Criteria Statistics
+A.1. N2C2 Criteria Statistics
 | Criterion | Met Count | Not Met Count |
@@
 | MI-6MOS | 26 | 262 |
-A.2 N2C2 Original Criteria
+A.2. N2C2 Original Criteria
 | Criteria | Definition |
@@
 | HBA1C | Any HbA1c value between 6.5 and 9.5%. |
-| CREATININE | Serum creatinine > upper limit of normal. |
-A.3 Redefined Criteria
+| CREATININE | Serum creatinine ¿ upper limit of normal. |
+A.3. Redefined N2C2 Criteria
 | Criteria | Definition |
@@
 | MI-6MOS | Myocardial infarction (MI) within the past 6 months. |
-A.4 MAKA-Augemented Criteria
+A.4. MAKAR-Augemented N2C2 Criteria
 | Criteria | Definition |
@@
 | Explanation: MI refers to heart muscle damage from blocked blood flow. Assess via medical history and records confirming the event within 6 months. |
+A.5. Sampled ClinicalTrial.gov Criteria
+| Inclusion Criteria | Exclusion Criteria |
+| 1.
+Locally advanced or metastatic breast cancer confirmed by histopathology.
+| 1.
+Patients known or suspected intolerance or hypersensitivity to main ingredient or any of the excipients of SNB-101.
+|
+| 2. Received 2 previous lines of anti-tumor treatment and developed resistance to standard treatment. | 2. Patients with bilateral invasive breast cancers. |
+| 3.
+Patients with locally advanced or metastatic advanced solid tumors confirmed by histology or cytology (including but not limited to triple-negative breast cancer, gastric cancer, colorectal cancer); Patients with 1 line of standard treatment failure (disease progression after treatment or intolerability of toxic side effects of treatment), or no standard treatment, or unable to receive standard treatment.
+| 3.
+Age 18 years old or 70 years old.
+|
+| 4.
+At least 1 measurable lesion per Response Evaluation Criteria in Solid Tumors version 1.1 and previously treated lesions with radiotherapy or focal therapy and no progression cannot be included as target lesion for assessment.
+| 4.
+Patients with standard metallic contraindications to CMR or estimated glomerular filtration rate 30 mL/min/1.73 m².
+|
+| 5.
+Histologically confirmed breast cancer with an invasive component measuring 20 mm and/or with morphologically confirmed spread to regional lymph nodes (stage cT2-cT4 with any cN, or cN1-cN3 with any cT).
+| 5.
+History of hypersensitivity or contraindication to TTF.
+|
+| 6.
+Known HER2-positive breast cancer defined as an IHC status of 3+.
+If IHC is 2+, a positive in situ hybridization (FISH, CISH, or SISH) test is required by local laboratory testing.
+| 6.
+Implanted pacemaker, defibrillator, or other electrical medical devices.
+|
+| 7.
+Patients with breast cancer planned to receive Anthracycline and Cyclophosphamide chemotherapy.
+| 7.
+Bilateral breast cancer (including multifocal breast cancer.
+All histological lesions were HER2 1+ or 2+, not detected by FISH amplification and HR positive according to ASCO guidelines).
+|
+| 8.
+Patients treated for hormone-dependent localized breast cancer requiring adjuvant hormonal therapy (HT).
+| 8.
+Received any form of anti-tumor therapy (chemotherapy, radiotherapy, molecular targeted therapy, endocrine therapy, etc.).
+|
+| 9. Patients treated with tamoxifen for a maximum of 1 to 3 years. | 9. Breast cancer without histopathological diagnosis. |
+| 10.
+Histologically confirmed ER+/HER2- early-stage resected invasive breast cancer at high or intermediate risk of recurrence, , based on clinical-pathological risk features, as defined in the protocol.
+| 10.
+Known personal history of ductal carcinoma in situ (DCIS) or invasive breast cancer.
+|
+| 11.
+Completed adequate (definitive) locoregional therapy (surgery with or without radiotherapy) for the primary breast tumour(s), with or without (neo)adjuvant chemotherapy.
+| 11.
+Prior systemic treatment for any malignancy.
+|
+| 12. Completed at least 2 years but no more than 5 years (+3 months) of adjuvant ET (+/- CDK4/6 inhibitor). | 12. Significant medical comorbidities as per investigator evaluation. |
+| 13. Patients receiving letrozole for more than two months. | 13. Patients with ongoing indications for the cardioprotective medication - ACE inhibitors, ARBs and/or beta-blockers |
+| 14. Postmenopausal hormone receptor-positive patients. | 14. Intermediate or high-grade ductal carcinoma in situ. |
+| 15. Patients with N0 disease or those who have undergone targeted axillary detection with a good response. | 15. Invasive carcinoma. |